# Efficient Bayesian analysis of kilonovae and GRB afterglows with fiesta

### v1

@@
11institutetext: Institut für Physik und Astronomie, Universität Potsdam, Haus 28, Karl-Liebknecht-Str. 24/25, 14476, Potsdam, Germany, 11email: hauke.koehn@uni-potsdam.de 22institutetext: Institute for Gravitational and Subatomic Physics (GRASP), Utrecht University, Princetonplein 1, 3584 CC Utrecht, The Netherlands, 22email: t.r.i.wouters@uu.nl 33institutetext: Nikhef, Science Park 105, 1098 XG Amsterdam, The Netherlands 44institutetext: Department of Physics and Earth Science, University of Ferrara, via Saragat 1, I-44122 Ferrara, Italy 55institutetext: INFN, Sezione di Ferrara, via Saragat 1, I-44122 Ferrara, Italy 66institutetext: INAF, Osservatorio Astronomico d’Abruzzo, via Mentore Maggini snc, 64100 Teramo, Italy 77institutetext: DTU Space, National Space Institute, Technical University of Denmark, Elektrovej 327/328, DK-2800 Kongens Lyngby, Denmark 88institutetext: Max Planck Institute for Gravitational Physics (Albert Einstein Institute), Am Mühlenberg 1, Potsdam 14476, Germany
Efficient Bayesian analysis of kilonovae and GRB afterglows with fiesta
[…]
Gamma-ray burst (GRB) afterglows and kilonovae (KNe) are electromagnetic transients that can accompany binary neutron star (BNS) mergers. Therefore, studying their emission processes is of general interest for constraining cosmological parameters or the behavior of ultra-dense matter. One common method to analyze electromagnetic data from BNS mergers is to sample a Bayesian posterior over the parameters of a physical model for the transient. However, sampling the posterior is computationally costly, and because of the many likelihood evaluations needed in this process, detailed models are too expensive to be used directly in inference. In the present article, we address this problem by introducing fiesta, a python package to train machine learning (ML) surrogates for GRB afterglow and kilonova models that can accelerate likelihood evaluations. Specifically, we introduce extensive ML surrogates for the state-of-the-art GRB afterglow models afterglowpy and pyblastafterglow, as well as a new surrogate for the KN emission as modeled by the possis code. Our surrogates enable evaluation of the lightcurve posterior within minutes. We also provide built-in posterior sampling capabilities in fiesta that rely on the flowMC package which scale efficiently to higher dimensions when adding up to tens of nuisance sampling parameters. Because of its use of the jax framework, fiesta also allows for GPU acceleration during both surrogate training and posterior sampling. We apply our framework to reanalyze AT2017gfo/GRB170817A and GRB211211A using our surrogates, thus employing the new pyblastafterglow model for the first time in Bayesian inference.
@@ 1 Introduction
Gamma-ray bursts (GRBs) and kilonovae (KNe) are types of electromagnetic transients that can originate from a binary neutron star (BNS) merger, as confirmed most prominently by the observation of AT2017gfo (Utsumi et al., 2017; Coulter et al., 2017; Andreoni et al., 2017; Shappee et al., 2017; Soares-Santos et al., 2017; Lipunov et al., 2017; Valenti et al., 2017; Díaz et al., 2017; Tanvir et al., 2017), and GRB170817A (Goldstein et al., 2017; Abbott et al., 2017b; Savchenko et al., 2017) in the wake of the gravitational wave (GW) event GW170817 (Abbott et al., 2017c, d). Additionally, several GRBs have been associated with potential subsequent KNe (Tanvir et al., 2013; Ascenzi et al., 2019; Troja et al., 2019b; Rastinejad et al., 2022; Troja et al., 2018; Levan et al., 2024; Yang et al., 2023; Levan et al., 2023; Stratta et al., 2025). To analyze these electromagnetic transients with Bayesian inference, a generative model is needed that links the source and observational parameters of the system (like ejecta masses, isotropic energy equivalent of the jet, observation angle) to the observed data. This model is then employed in a sampling procedure to obtain a posterior distribution.
@@
To model the KN emission, one has to determine the quasi-thermal radiation that is emitted by the BNS ejecta heated from the radioactive decay of the synthesized nuclei. This process has been investigated through various methods and approaches in the literature, see e.g. Kasen et al. (2017); Villar et al. (2017); Kawaguchi et al. (2018); Metzger (2020); Breschi et al. (2021); Wollaeger et al. (2021); Curtis et al. (2022); Nicholl et al. (2021); Bulla (2023). The GRB prompt emission, on the other hand, takes place on the scales of seconds to minutes and is observed as bursts of highly energetic gamma and X-ray radiation. Its emission mechanism is largely uncertain, and thus the associated data are typically not taken into account for Bayesian inference of BNS mergers. However, the prompt emission is followed by an afterglow with broadband radiation spanning from the radio to the X-ray, which lasts on a timescale of minutes to years (Miceli & Nava, 2022). Unlike the GRB prompt emission, the afterglow physics is comparatively well understood and can thus be used to infer properties of the jet and the progenitor. Specifically, the afterglow emission arises from the interaction of the GRB jet with the surrounding cold interstellar medium, and various afterglow models are available in the literature, e.g., in van Eerten et al. (2012); Ryan et al. (2015); Lamb et al. (2018); Ryan et al. (2020); Zhang et al. (2021); Pellouin & Daigne (2024); Wang et al. (2024); Nedora et al. (2024); Wang et al. (2025).
@@
Such models have been successfully applied in joint Bayesian analysis of GW170817, its KN, and the GRB afterglow to place constraints on the equation of state (EOS) for neutron star (NS) matter (Radice et al., 2018b; Dietrich et al., 2020; Raaijmakers et al., 2021; Nicholl et al., 2021; Pang et al., 2023; Güven et al., 2020; Annala et al., 2022; Breschi et al., 2024; Koehn et al., 2025) or to determine the Hubble constant (Abbott et al., 2017a; Hotokezaka et al., 2019; Dietrich et al., 2020; Mukherjee et al., 2021; Wang & Giannios, 2021; Gianfagna et al., 2024). However, these joint multi-messenger inferences pose certain challenges, as exploring their high-dimensional parameter space is computationally expensive. Additionally, evaluating the likelihood at a given parameter point requires calculating the expected emission for these given parameters. Therefore, the direct, physical calculation of the emission on-the-fly is only viable with computationally cheap, semi-analytical models. However, even relatively efficient models can become prohibitive, when considering multi-messenger inferences of BNS mergers signals, where GW and electromagnetic signals are analyzed jointly and due to the increased dimensionality, many more likelihood evaluations are needed. Thus, if one wants to use a more expensive, involved model in Bayesian multi-messenger inference, one has to build accurate surrogate models, typically involving machine learning (ML) techniques, in order to evaluate the likelihood with sufficient speed. Furthermore, joint multi-messenger BNS inferences also provide motivation for the GPU compatible lightcurve surrogates, since the analysis of BNS GW data can be significantly accelerated through GPU hardware (Wysocki et al., 2019; Wouters et al., 2024; Hu et al., 2025; Dax et al., 2025).
@@
In the present article, we introduce fiesta, a jax-based python package for training ML surrogates of KN and GRB afterglow models and for the Bayesian analysis of photometric transient lightcurves. With fiesta, we provide extensive ML surrogates that effectively replace the costly evaluation of the physical base model, enabling rapid evaluation of the expected lightcurve given some parameter values. Specifically, we present surrogates that are built upon GRB afterglow models from the popular afterglowpy model (Ryan et al., 2020) and the recently developed pyblastafterglow (Nedora et al., 2024). Additionally, we introduce a new KN surrogate based on the 3D Monte-Carlo radiation transport code possis (Bulla, 2019, 2023). Contrary to many previous works, our surrogates are not trained on the integrated photometric flux in a specific bandpass, but predict the entire spectral flux density, providing maximal flexibility. Given photometric transient data, fiesta’s surrogates can be used in stochastic samplers to achieve fast evaluation of the likelihood, which opens the door for fast transient analysis. Such analyses can be conducted using the established inference framework nmma (Pang et al., 2023) with which our surrogates are compatible. In addition, fiesta also contains its own sampling implementation that relies on the flowMC package (Wong et al., 2023a) to evaluate the posterior MCMC chain with normalizing flows and the Metropolis-adjusted Langevin algorithm (MALA). These advanced sampling techniques reduce the number of likelihood evaluations needed and thus sample the posterior more efficiently, therefore improving the scaling of fiesta as we consider additional nuisance parameters to account for systematic uncertainties. Because fiesta uses the jax framework, surrogate training as well as posterior sampling can be GPU-accelerated.
@@
The approach implemented in fiesta enables sampling the full lightcurve posterior within minutes, which previously, depending on the base model, would have either taken several hours to days, or been prohibitively expensive from the start. Therefore, the fiesta surrogates make previously intractable models available for Bayesian inference, which allows us to present the first Bayesian analyses of GRB afterglows with pyblastafterglow. The fiesta code together with the surrogates is publicly available [^1] https://github.com/nuclear-multimessenger-astronomy/fiestaEM and all the data used in the present article can be accessed as well [^2] https://github.com/nuclear-multimessenger-astronomy/paper_fiesta/.
@@
The present article is organized as follows: In Sec. 2, we briefly review Bayesian inference of photometric transient data. We then continue in Sec. 3 discussing machine learning approaches to create surrogate models for the KN and GRB afterglow emission and present our flagship surrogates for afterglowpy, pyblastafterglow, and possis. From there, we verify that our surrogates are able to accurately recover the posterior and discuss their performance in Sec. 4. In Sec. 5, we apply fiesta to analyze the data from AT2017gfo/GRB170817A and GRB211211A before concluding in Sec. 6.
@@ 2 Bayesian inference of transients
Given photometric light curve data $d$, the goal is to find the posterior $P(\@vec{\theta}|d)$, where ${\theta}$ denotes the parameters of the model that describes the physical emission process. The posterior can be obtained by using the Bayes’ theorem:
@@
where $\mathcal{L}(\@vec{\theta}|d)$ is called the likelihood function, $\pi(\@vec{\theta})$ the prior distribution, and $Z$ the Bayesian evidence. The latter is an important quantity for model selection (e.g., jet geometries in the context of GRBs). Since the posterior is often analytically intractable, stochastic sampling methods are used. One technique is nested sampling (Skilling, 2004, 2006), which computes the Bayesian evidence (i.e., the normalization constant of the posterior distribution, see Eq. (1)), from which posterior samples can be obtained as a byproduct. Alternatively, Markov-chain Monte Carlo (MCMC) methods (Neal, 2011), which directly generate samples from the posterior. Direct MCMC methods do not provide an estimate of the evidence, unless they are supplemented with additional techniques such as parallel tempering (Marinari & Parisi, 1992) or the learned harmonic mean estimator (Newton & Raftery, 1994; McEwen et al., 2023; Polanska et al., 2024). However, we do not consider these approaches in the present article.
In the context of photometric lightcurve analyses, we denote the observed data $d$ as a time series of magnitudes $\{m(t_{j})|j=1,2,3\dots\}$, and the corresponding predictions of the model by $m^{\star}(t_{j},\@vec{\!\theta}\,)$. The corresponding likelihood function can be written as
@@
Here, $\sigma(t_{j})$ stands for the measurement uncertainty of each data point, and $\sigma_{\text{sys}}(t_{j})$ is the model systematic uncertainty that accounts for both the surrogate error and the systematic offset caused by simplifying assumptions in the physical base model. When $\sigma_{\text{sys}}(t)$ is sampled freely, either as single inference parameter or determined by several parameters, the second term in Eq. (2) serves to offset the decrease of the first term and forces the sampler to pick the smallest systematic uncertainty that still matches the data.
Existing inference frameworks such as nmma (Pang et al., 2023), bajes (Breschi et al., 2024), redback (Sarin et al., 2024), or MOSFiT (Guillochon et al., 2018), evaluate a model or a surrogate to determine the value of Eq. (2) and use either nested sampling or MCMC methods to sample the posterior. fiesta provides surrogates that can be used with nmma, and potentially other inference frameworks, but also contains its own sampling implementation based on the flowMC sampler (Wong et al., 2023a). While nmma relies on nested sampling through the PyMultinest (Feroz et al., 2009) and dynesty (Speagle, 2020) package, fiesta has sampling functionalities that utilize flowMC. The latter is an MCMC sampler enhanced by gradient-based sampling (in particular, the Metropolis-adjusted Langevin algorithm (Grenander & Miller, 1994)) and normalizing flows, which are a class of generative ML methods that act as neural density estimators (Rezende & Mohamed, 2015; Papamakarios et al., 2021; Kobyzev et al., 2020). The flows are trained on the fly from the MCMC chains and subsequently used as proposal distributions in an adaptive MCMC algorithm (Gabrié et al., 2022).
In both fiesta and nmma, the systematic error $\sigma_{\text{sys}}(t)$ in Eq. (2) can either be fixed to a constant value or be sampled freely from a prior. Moreover, fiesta and nmma support time- and filter-dependent systematic uncertainties by sampling the nuisance parameters $\sigma_{\text{sys}}^{(k)}$ at specific time nodes $t_{k}$. The systematic error for a filter $f$ at a specific data point $t_{j}$ is then determined through linear interpolation
$\sigma_{\text{sys}}(t_{j},f)=\sigma_{\text{sys}}^{(k)}(f)+\frac{t_{j}-t_{k}}{t_{k+1}-t_{k}}(\sigma_{\text{sys}}^{(k+1)}(f)-\sigma_{\text{sys}}^{(k)}(f))$ (3)
and this value is then used in Eq. (2). The nuisance parameters $\sigma_{\text{sys}}^{(k)}(f)$ can be sampled for different filters separately. This implementation for a data-driven inference of the systematic uncertainty essentially follows Jhawar et al. (2025).

### v2

# Efficient Bayesian analysis of kilonovae and gamma ray burst afterglows with fiesta

@@
11institutetext: Institut für Physik und Astronomie, Universität Potsdam, Haus 28, Karl-Liebknecht-Str. 24/25, 14476, Potsdam, Germany, 11email: hauke.koehn@uni-potsdam.de 22institutetext: Institute for Gravitational and Subatomic Physics (GRASP), Utrecht University, Princetonplein 1, 3584 $C$ Utrecht, The Netherlands, 22email: t.r.i.wouters@$u$.nl 33institutetext: Nikhef, Science Park 105, 1098 XG Amsterdam, The Netherlands 44institutetext: Department of Physics and Earth Science, University of Ferrara, via Saragat 1, I-44122 Ferrara, Italy 55institutetext: INFN, Sezione di Ferrara, via Saragat 1, I-44122 Ferrara, Italy 66institutetext: INAF, Osservatorio Astronomico d’Abruzzo, via Mentore Maggini snc, 64100 Teramo, Italy 77institutetext: DTU Space, National Space Institute, Technical University of Denmark, Elektrovej 327/328, DK-2800 Kongens Lyngby, Denmark 88institutetext: Max Planck Institute for Gravitational Physics (Albert Einstein Institute), Am Mühlenberg 1, Potsdam 14476, Germany
Efficient Bayesian analysis of kilonovae and gamma ray burst afterglows with fiesta
H. Koehn    T. Wouters    P. T. H. Pang    M. Bulla    H. Rose    H. Wichern    T. Dietrich
(Received 28 July 2025 / Accepted 20 October 2025)
Gamma-ray burst (GRB) afterglows and kilonovae (KNe) are electromagnetic transients that can accompany binary neutron star (BNS) mergers. Therefore, studying their emission processes is of general interest for constraining cosmological parameters or the behavior of ultra-dense matter. One common method to analyze electromagnetic data from BNS mergers is to sample a Bayesian posterior over the parameters of a physical model for the transient. However, sampling the posterior is computationally costly and because of the many likelihood evaluations required in this process, detailed models are too expensive to be used directly in Bayesian inference. In this paper, we address the problem by introducing fiesta, a python package to train machine learning (ML) surrogates for GRB afterglow and kilonova models that have the capacity to accelerate likelihood evaluations. Specifically, we introduce extensive ML surrogates for the state-of-the-art GRB afterglow models afterglowpy and pyblastafterglow, along with a new surrogate for KN emission based on the possis code. Our surrogates enable evaluation of the light-curve posterior within minutes. We also provide built-in posterior sampling capabilities in fiesta that rely on the flowMC package, which efficiently scale to higher dimensions when adding up to tens of nuisance sampling parameters. Because of its use of the jax framework, fiesta also allows for GPU acceleration during both surrogate training and posterior sampling. We applied our framework to reanalyze AT2017gfo/GRB170817A and GRB211211A with our surrogates, thus employing the new pyblastafterglow model for the first time in Bayesian inference.
@@ 1 Introduction
Gamma-ray bursts (GRBs) and kilonovae (KNe) are types of electromagnetic transients that can originate from a binary neutron star (BNS) merger, as confirmed most prominently by the observation of AT2017gfo (J-GEM:2017tyx; Coulter:2017wya; Andreoni:2017ppd; Shappee:2017zly; DES:2017kbs; Lipunov:2017dwd; Valenti:2017ngx; TOROS:2017pqe; Tanvir:2017pws) and GRB170817A (Goldstein:2017mmi; LIGOScientific:2017zic; Savchenko:2017ffs) in the wake of the gravitational wave (GW) event GW170817 (LIGOScientific:2017vwq; LIGOScientific:2017ync). Additionally, several GRBs have been associated with potential subsequent KNe (Tanvir:2013pia; Ascenzi:2018mbh; Troja:2019ccb; Rastinejad:2022zbg; Troja:2018ybt; JWST:2023jqa; Yang:2023mqt; Levan:2023qdh; Stratta:2024kbs). To analyze these electromagnetic transients with Bayesian inference, a physical model that links the source and observational parameters of the system (e.g., ejecta masses, isotropic energy equivalent of the jet, observation angle) to the observed data is needed. This model can then be employed in a sampling procedure to obtain a posterior distribution.
@@
The KN emission arises from quasi-thermal radiation produced by the BNS ejecta heated from the radioactive decay of nuclei synthesized through the r-process (Thielemann:2011) and is observed on the timescale of days after the merger. This process has been investigated through various methods and approaches in the literature (e.g., Kasen:2017sxr; Villar:2017wcc; Kawaguchi:2018ptg; Metzger:2019zeh; Breschi:2021tbm; Wollaeger:2021qgf; Curtis:2021guz; Nicholl:2021rcr; Bulla:2022mwo). The GRB prompt emission, on the other hand, takes place on scales of seconds to minutes and is observed as bursts of highly energetic gamma and X-ray radiation. Its emission mechanism is largely uncertain and, thus, the associated data are typically not taken into account for Bayesian inference of BNS mergers. However, the prompt emission is followed by an afterglow of broadband radiation spanning from radio to $\gamma$-ray frequencies, observable on timescales ranging from minutes to years (Miceli:2022efx). Unlike the GRB prompt emission, the afterglow physics are comparatively well understood and can thus be used to infer properties of the jet and the progenitor. Specifically, the afterglow emission arises from the interaction of the GRB jet with the surrounding cold interstellar medium and various afterglow models are available in the literature (e.g., vanEerten:2011yn; Ryan:2014nea; Lamb:2018ohw; Ryan:2019fhz; Zhang:2020tem; Pellouin:2024gqj; Wang:2024wbt; Nedora:2024vrv; Wang:2025ccz).
@@
Such models have been successfully applied in joint Bayesian analysis of GW170817, its KN, and the GRB afterglow to place constraints on the equation of state (EOS) for neutron star (NS) matter (Radice:2017lry; Dietrich:2020efo; Raaijmakers:2021uju; Nicholl:2021rcr; Pang:2022rzc; Guven:2020dok; Annala:2021gom; Breschi:2024qlc; Koehn:2024set) or to determine the Hubble constant (LIGOScientific:2017adf; Hotokezaka:2018dfi; Dietrich:2020efo; Mukherjee:2019qmm; Wang:2020vgr; Gianfagna:2023cgk). However, these joint multi-messenger inferences pose certain computational challenges, because exploring their high-dimensional parameter space requires many likelihood evaluations. Therefore, a single likelihood evaluation needs to be as cheap as possible in order to keep the total sampling time manageable. However, to evaluate the likelihood function at a given parameter point, we have to determine the expected emission for these given parameters from the physical model. Due to the cost limit on the likelihood function, any direct, physical calculation of the emission on-the-fly is only viable with computationally cheap, semi-analytical models. Yet, even relatively efficient models can become prohibitive, when considering multi-messenger inferences of BNS mergers signals, where GW and electromagnetic signals are analyzed jointly in a large parameter space. Accordingly, applying more expensive and involved models in multi-messenger inference necessitates the development of accurate surrogate models, often through machine learning (ML) techniques to enable likelihood evaluation at sufficient speeds (Pang:2022rzc; Almualla:2021znj; Kedia:2022onl; Ristic:2021ksz; Boersma:2022qtg). Furthermore, joint multi-messenger BNS inferences also provide motivation for the GPU compatible light-curve surrogates, since the analysis of BNS GW data can be significantly accelerated on GPU hardware (Wysocki:2019grj; Wouters:2024oxj; Hu:2024lrj; Dax:2024mcn).
@@
In the present article, we introduce fiesta, a jax-based python package for training ML surrogates of KN and GRB afterglow models and for the Bayesian analysis of photometric transient light curves. With fiesta, we provide extensive ML surrogates that effectively replace the costly evaluation of the physical base model, enabling rapid prediction of the expected light curve given the model’s parameters. Specifically, we present surrogates that are built upon GRB afterglow models from the popular afterglowpy model (Ryan:2019fhz) and the recently developed pyblastafterglow (Nedora:2024vrv). Additionally, we introduce a new KN surrogate based on the 3D Monte Carlo radiation transport code possis (Bulla:2019muo; Bulla:2022mwo). In contrast to many previous works, our surrogates have not been trained on the magnitudes in a specific photometric passband; instead, they predict the entire spectral flux density, providing maximal flexibility. Given photometric transient data, fiesta’s surrogates can be used in stochastic samplers to achieve fast evaluation of the likelihood, which opens the door for swift transient analysis. Such analyses can be conducted using the established inference framework nmma (Pang:2022rzc), which our surrogates are compatible with. In addition, fiesta also contains its own sampling implementation that relies on the flowMC package (Wong:2022xvh) to generate the posterior Markov chain Monte Carlo (MCMC) chain with normalizing flows and the Metropolis-adjusted Langevin algorithm (MALA). These advanced sampling techniques reduce the number of likelihood evaluations needed and thus sample the posterior more efficiently, thereby improving the scaling of fiesta as we consider additional nuisance parameters to account for systematic uncertainties. Because fiesta uses the jax framework, surrogate training as well as posterior sampling can be GPU-accelerated.
@@
The approach implemented in fiesta enables sampling the full light-curve posterior within minutes, which, depending on the base model, would previously have either taken several hours to days or proven prohibitively expensive from the start. Therefore, the fiesta surrogates make previously intractable models available for Bayesian inference, which allows us to present the first Bayesian analyses of GRB afterglows with pyblastafterglow. The fiesta code together with the surrogates is publicly available [^1] https://github.com/nuclear-multimessenger-astronomy/fiestaEM and all the data used in the present article can be accessed as well [^2] https://github.com/nuclear-multimessenger-astronomy/paper_fiesta/.
@@
The present article is organized as follows: In Sect. 2, we briefly review Bayesian inference of photometric transient data. We then continue in Sect. 3 discussing machine learning approaches to create surrogate models for the KN and GRB afterglow emission and present our flagship surrogates for afterglowpy, pyblastafterglow, and possis. From there, we verify that our surrogates are able to accurately recover the posterior and discuss their performance in Sect. 4. In Sect. 5, we apply fiesta to analyze the data from AT2017gfo/GRB170817A and GRB211211A before concluding in Sect. 6. Throughout, we use the AB magnitude system to convert interchangeably between flux densities, $F_{\nu}$, and magnitudes,
$m=-2.5\ \log_{10}\left(\frac{\int F_{\nu}\frac{e(\nu)}{h\nu}d\nu}{\int 3631\penalty 10000\ \text{Jy}\ \frac{e(\nu)}{h\nu}d\nu}\right),$ (1)
where $e(\nu)$ is the detector response function. In this way, even X-ray or radio flux density measurements can be expressed as magnitudes within fiesta.
@@ 2 Bayesian inference of transients
Given some photometric light curve data $d$, the goal is to find the posterior $P(\@vec{\theta}|d)$, where $\@vec{\theta}$ denotes the parameters of the model that describes the physical emission process. The posterior can be obtained by using Bayes’ theorem,
@@
where $\mathcal{L}(\@vec{\theta}|d)$ is called the likelihood function, $\pi(\@vec{\theta})$ the prior distribution, and $Z$ is the Bayesian evidence. The latter is an important quantity for model selection (e.g., jet geometries in the context of GRBs). Since the posterior is often analytically intractable, stochastic sampling methods are used. One technique is nested sampling (Skilling:2004pqw; Skilling:2006gxv), which computes the Bayesian evidence (i.e., the normalization constant of the posterior distribution, see Eq. (2)), from which posterior samples can be obtained as a byproduct. Alternatively, we can look to MCMC methods (Neal:2011mrf), which directly generate samples from the posterior. Direct MCMC methods can only provide an estimate of the evidence if they are supplemented with additional techniques such as parallel tempering (Marinari:1992qd) or the learned harmonic mean estimator (Newton:1994; mcewen2023machinelearningassistedbayesian; Polanska:2024arc). However, we do not consider these approaches in the present article.
In the context of photometric light curve analyses, we denote the observed data, $d$, as a time series of magnitudes $\{m(t_{j})|j=1,2,3\dots\}$, and the corresponding predictions of the model as $m^{\star}(t_{j},\@vec{\!\theta}\,)$. The data are taken with some measurement uncertainty, $\sigma(t_{j})$. In fiesta, we assume that this uncertainty is Gaussian and hence the corresponding likelihood function can be written as
@@
Here, $\sigma_{\text{sys}}(t_{j})$ is the model systematic uncertainty that accounts for both the surrogate error and the systematic offset caused by simplifying assumptions in the physical base model. Moreover, detection limits can also be incorporated into fiesta’s likelihood. For every detection limit $m^{\blacktriangledown}(t_{j})$, the likelihood in Eq. (3) is simply multiplied by
$\mathcal{L}$ $(\@vec{\theta}|d)=\mathcal{L}(\@vec{\theta}|d)\times\int^{\infty}_{m^{\blacktriangledown}(t_{j})}\frac{\exp\left(-\frac{1}{2}\bigl(\frac{x-m^{\star}(t_{j},\@vec{\theta})}{\sigma_{\rm sys}(t_{j})}\bigr)^{2}\right)}{\sqrt{2\pi\sigma_{\rm sys}(t_{j})^{2}}}\ dx\ .$ (4)
We note that depending on the detector and measurement, other likelihood statistics might be more appropriate. For instance, for low-flux X-ray data, the Poisson distribution is generally better suited to describe the measurement uncertainty (Ryan:2023pzk; Humphrey:2008xz). In fiesta, we stick to a Gaussian likelihood and assume equal upper and lower error bars.
Existing inference frameworks such as nmma (Pang:2022rzc), bajes (Breschi:2024qlc), redback (Sarin:2023khf), or MOSFiT (Guillochon:2017bmg) evaluate a model or a surrogate to determine the value of Eq. (3). They either use nested sampling or MCMC methods to sample the posterior. fiesta provides surrogates that can be used with nmma and potentially other inference frameworks, but also contains its own sampling implementation based on the flowMC sampler (Wong:2022xvh). While nmma relies on nested sampling through the PyMultinest (Feroz:2008xx) and dynesty (Speagle:2019ivv) packages, fiesta has sampling functionalities that utilize flowMC. The latter is an MCMC sampler enhanced by gradient-based sampling (in particular, the Metropolis-adjusted Langevin algorithm (grenander1994representations)) and normalizing flows, which are a class of generative ML methods that act as neural density estimators (Rezende:2015ocs; Papamakarios:2019fms; kobyzev2020normalizing). The flows are trained on the fly from the MCMC chains and subsequently used as proposal distributions in an adaptive MCMC algorithm (Gabrie:2021tlu).
In both fiesta and nmma, the systematic error $\sigma_{\text{sys}}(t)$ in Eq. (3) can either be fixed to a constant value or be sampled freely from a prior. Moreover, fiesta and nmma support time- and filter-dependent systematic uncertainties by sampling the nuisance parameters $\sigma_{\text{sys}}^{(k)}$ at specific time nodes $t_{k}$. The systematic error for a filter, $f$, at a specific data point, $t_{j}$, is then determined via a linear interpolation,
$\sigma_{\text{sys}}(t_{j},f)=\sigma_{\text{sys}}^{(k)}(f)+\frac{t_{j}-t_{k}}{t_{k+1}-t_{k}}(\sigma_{\text{sys}}^{(k+1)}(f)-\sigma_{\text{sys}}^{(k)}(f)),$ (5)
which is then placed into Eq. (3). The nuisance parameters, $\sigma_{\text{sys}}^{(k)}(f)$, can be sampled separately for different filters. This implementation for a data-driven inference of the systematic uncertainty essentially follows Jhawar:2024ezm.